OpenStudy (anonymous):

Jon makes a mix of salad that costs $15 using as much apples that cost $4/kg as oranges that cost $6/kg. How much of each fruit is in the salad?

OpenStudy (aihberkhan):

Okay. 4x+6y = 15 x = y 4x+6x = 15 x = 1.5 kg y = 1.5 kg

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @dianneca

OpenStudy (whpalmer4):

Another way of thinking about this problem: we know that we have equal quantities of the two fruits. One costs $4/kg, the other costs $6/kg. If we have 1 kg of each, the cost of that will be $4 + $6 = $10. How many of those 2 kg batches do we need to make the cost be $15? $15/$10 = 1.5, so we need 1.5 of those batches. 1.5 * 1 kg = 1.5 kg, so we have 1.5 kg of each fruit. Checking our work, if we multiply the amount of each fruit by its price and add them up, we should get $15 \[1.5\text{ kg}*{\$4/\text{kg}} + 1.5\text{ kg}*{\$6/\text{kg}} = \$6 + \$9 = \$15\checkmark \]